Exercise 11-12 (Video) Byrd Company produces one product, a putter called GO-Putter. Byrd uses a standard cost system and determines that it should take one hour of direct labor to produce one GO-Putter. The normal production capacity for this putter is 100,000 units per year. The total budgeted overhead at normal capacity is $850,000 comprised of $300,000 of variable costs and $550,000 of fixed costs. Byrd applies overhead on the basis of direct labor hours. During the current year, Byrd produced 81,800 putters, worked 97,800 direct labor hours, and incurred variable overhead costs of $173,825 and fixed overhead costs of $642,300. Compute the predetermined variable overhead rate and the predetermined fixed overhead rate. (Round answers to 2 decimal places, e.g. 2.75.) Variable Fixed Predetermined Overhead Rate $ $ LINK TO TEXT Compute the applied overhead for Byrd for the year. Overhead Applied $ LINK TO TEXT Compute the total overhead variance. Total Overhead Variance $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

Budgeted hours = 100,000*1 = 100,000 hours Predetermined variable overhead rate = $300,000/100,000 = $3 per DLH Predetermined